OpenStudy (anonymous):

need help.. describe this sets: {2,3,5,7,11} {-3,-2,-1,0,1,2} {0,1,4,9,16,25} {-8,-1,0,1,8,27} {...,-3,-2,-1,0,1,2,...}

OpenStudy (anonymous):

Last one is\[ \mathbb{Z} \]

OpenStudy (anonymous):

First is the set of the 5 first prime number.

OpenStudy (anonymous):

what do you mean by last one is a Z

OpenStudy (anonymous):

Third is the suit of square : 1², 2², 3²; 4² ... Last one if the set of integer : http://en.wikipedia.org/wiki/Integer

OpenStudy (anonymous):

Z is the set of all integers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think the fourth one is the set of all perfect cubes between -8 and 27

OpenStudy (anonymous):

how would you describe that in .... form: like x "is a subset of " (-8,0) (0, infinity] could that be annoted like that? (i know that is incorrect already from the ))

OpenStudy (anonymous):

jimmyrep how about the rest?

OpenStudy (anonymous):

paulzord : Set builder ? \begin{Bmatrix} x / x = y^3 , y \in \left \{-8,...,27 \right \} \end{Bmatrix}

OpenStudy (anonymous):

yep - thats right william

OpenStudy (anonymous):

how about the second one ? thats a subset of Z, right?

OpenStudy (anonymous):

to be honest i'm a little rusty in set theory

OpenStudy (anonymous):

Would say somethig like that : \[\begin{Bmatrix} x \: | \: x \in [-3 ; 2] \end{Bmatrix} \] but seems a bit obvious. Integer modulo n perhaps ?

OpenStudy (anonymous):

Or, for integer only: \[\begin{Bmatrix} x \: | \: x \in |[-3 ; 2]| \end{Bmatrix} \]